# Learning normalized image densities via dual score matching

### v1

@@ 4 Discussion
We have developed a novel framework for estimating the log probability (energy) of a distribution of images from observed samples. The estimation method is simple and robust, and converges to a stable solution with relatively small amounts of data (in our examples, $100,000$ images of size $64\times 64$ suffices). The framework leverages the tremendous power of generative diffusion models, relying on networks trained to estimate the score of the data distribution by minimizing a denoising objective. We augment this with a secondary objective that ensures consistency of the model across noise levels, and an architecture constructed from an existing denoiser so as to preserve its inductive biases. We validate our model by verifying that it achieves denoising performance equal to or surpassing the original denoiser, and NLL values comparable to the state of the art.
We have investigated the geometrical properties of the learned energy model. Notably, we observe a lack of concentration of measure, with log probability values varying over a wide range and following a Gumbel distribution. As a limit distribution for the maximum, rather than the sum, of i.i.d. random variables, we expect that this might arise in a broad class of high-dimensional probability distributions. We also introduced a novel image- and scale-dependent effective dimensionality measure which unifies two separate points of view adopted in the literature. We demonstrate that different image neighborhoods display both high and low dimensionality over a wide range of scales. These results challenge simple interpretations of the manifold hypothesis. Furthermore, the estimated dimensionalities remain too high to explain how the curse of dimensionality may be lifted. This indicates that the energy landscape of natural images must have additional geometrical regularity.
Limitations
We mention two main limitations of our work. First, our energy model has a roughly doubled training time compared to the base score network, due to the use of double back-propagation. We believe this can be improved through architecture improvements or specialized auto-differentiation functions. Another limitation of our approach is that we cannot provide any theoretical guarantees that minimizing our objective leads to a quantitative approximation of the energy (see Section A.3).
@@ A.1 On comparison of NLL values
Finally, a third point is that depending on the method, the NLL should be interpreted differently. Classically, $p_{\theta}$ is a normalized probability distribution, typically obtained through a change-of-variable formula (as in flow-based models), so that up to an unknown additive constant (the entropy of the data), the NLL directly evaluates the KL divergence $\operatorname{KL}\mathopen{}\mathclose{{}\left(p\,\middle\|\,p_{\theta}}\right)$. For some other models (such as VAEs), the NLL is not directly tied to a probabilistic model and is instead evaluated through a variational lower bound, leading to upper bounds on $-\log p(x)$ for each $x$. The NLL however remains an upper bound on the entropy of the data. The MSE-based formulas of Kong et al. (2023) also fall in this category when using a (necessarily suboptimal) network denoiser. Lastly, approximately normalized energy-based models such as ours or CDM (Yadin et al., 2024) compute estimates of $-\log p(x)$ that are neither lower nor upper bounds, so that NLL can only be interpreted as an estimate of the entropy of the data.
@@
When evaluating previously published NLL results, we found that these details were often not provided, and we thus had to make assumptions in compiling Table 2. The “anti-aliasing” column refers to the version of ImageNet64 used, we assumed that articles that cite Van den Oord et al. (2016) for the dataset made use of the aliased version. We assume no data augmentation if none is mentioned (for TarFlow (Zhai et al., 2024), the authors mention performing center crops of the data, but an inspection of their code indicates that it has been disabled for ImageNet64). The “discreteness” column refers to the nature of the probability model and the potential conversion from continuous to discrete probabilities (“discrete” for discrete probability models, “continuous” for an additive shift of $8$ bits, and “uniform” for uniform dequantization by adding uniform noise). We assume a continuous model if no dequantization is mentioned (FM (Lipman et al., 2023) and NFDM (Bartosh et al., 2024) use a slightly different notion of uniform dequantization which improves NLL by 0.04 bits/dimension). The “type” column refers to the nature of the LLM estimate (“normalized” for an exactly normalized probability model coming from a flow-based model, “upper bound” for variational lower bounds on log probability, and “estimate” for other approaches). Finally, we indicate those methods that use a single neural function evaluation (NFE) to compute log probability in the “single NFE” column.
@@ A.2 Background on density computation in diffusion models
$-\frac{\mathrm{d}x_{t}}{\mathrm{d}t}=-\frac{1}{2}\nabla_{y}U_{\theta}(x_{t},t),$ (13)
@@
$-\log p_{\rm ODE}(x)=\frac{{\mathopen{}\mathclose{{}\left\lVert x_{t_{\max}}}\right\rVert}^{2}}{2t_{\max}}+\frac{d}{2}\log\mathopen{}\mathclose{{}\left(2\pi t_{\max}}\right)-\frac{1}{2}\int_{0}^{t_{\max}}\Delta_{y}U_{\theta}(x_{t},t)\mathrm{d}t.$ (14)
Note that eq. 14 is also valid for arbitrary test points $x$, in which case the ODE (13) needs to be solved forward in time from $x_{0}=x$ at $t=0$ to $t=t_{\max}$. Equation 14 requires estimating the divergence of the score (the Laplacian of the energy), which is typically approximated with the Hutchinson trace estimator (Hutchinson, 1989).
@@
$-\log p(x)=\mathbb{E}_{y}\mathopen{}\mathclose{{}\left[U(y,t_{\max})\,\middle|\,x}\right]-\int_{0}^{t_{\max}}\mathopen{}\mathclose{{}\left(td-\mathbb{E}_{y}\mathopen{}\mathclose{{}\left[{\mathopen{}\mathclose{{}\left\lVert x-\mathbb{E}_{x}\mathopen{}\mathclose{{}\left[x\,\middle|\,y}\right]}\right\rVert}^{2}\,\middle|\,x}\right]}\right)\frac{\mathrm{d}t}{2t^{2}}.$ (15)
The first term is the effective energy at $t=t_{\max}$, which is equivalent to $\frac{d}{2}\log\mathopen{}\mathclose{{}\left(2\pi\mathrm{e}t_{\max}}\right)$ as $t_{\max}\to\infty$. The second term features the optimal mean squared error over noise levels $t\in[0,t_{\max}]$. Note that the integrand can be rewritten as $(d-d_{\rm eff}(x,t))\mathrm{d}\mathopen{}\mathclose{{}\left(\frac{1}{2}\log t}\right)$, and eq. 15 can thus be derived by integrating the two equivalent definitions of effective dimensionality (11) and (12) (see Section B.3). Equation 15 naturally leads to an upper-bound on the negative log probability of the data when replacing the optimal denoiser $\mathbb{E}_{x}\mathopen{}\mathclose{{}\left[x\,\middle|\,y,t}\right]$ with the denoiser derived from the Miyasawa-Tweedie expression, $y-t\nabla_{y}U_{\theta}(y,t)$:
$-\log p_{\rm MSE}(x)=\frac{d}{2}\log\mathopen{}\mathclose{{}\left(2\pi\mathrm{e}t_{\max}}\right)-\int_{0}^{t_{\max}}\mathopen{}\mathclose{{}\left(td-\mathbb{E}_{y}\mathopen{}\mathclose{{}\left[{\mathopen{}\mathclose{{}\left\lVert x-y+t\nabla_{y}U_{\theta}(y,t)}\right\rVert}^{2}\,\middle|\,x}\right]}\right)\frac{\mathrm{d}t}{2t^{2}}.$ (16)
@@
$-\log p(x)=U(x_{t_{\max}},t_{\max})-\int_{0}^{t_{\max}}\mathopen{}\mathclose{{}\left(\mathopen{}\mathclose{{}\left\langle\nabla_{y}U(x_{t},t),\mathrm{d}x_{t}}\right\rangle-\frac{1}{2}{\mathopen{}\mathclose{{}\left\lVert\nabla_{y}U(x_{t},t)}\right\rVert}^{2}\mathrm{d}t}\right).$ (17)
@@
$-\log p_{\rm SDE}(x)=\frac{{\mathopen{}\mathclose{{}\left\lVert x_{t_{\max}}}\right\rVert}^{2}}{2t_{\max}}+\frac{d}{2}\log\mathopen{}\mathclose{{}\left(2\pi t_{\max}}\right)-\int_{0}^{t_{\max}}\mathopen{}\mathclose{{}\left(\mathopen{}\mathclose{{}\left\langle\nabla_{y}U_{\theta}(x_{t},t),\mathrm{d}x_{t}}\right\rangle-\frac{1}{2}{\mathopen{}\mathclose{{}\left\lVert\nabla_{y}U_{\theta}(x_{t},t)}\right\rVert}^{2}\mathrm{d}t}\right).$ (18)
Karczewski et al. (2025) show that averaging over SDE trajectories started at the same $x_{0}=x$ leads to an upper bound on the NLL, in fact recovering the one in eq. 16.
In summary, the ODE gives deterministic probabilities exactly corresponding to the corresponding generative model (eq. 14), while for the SDE one can choose between an denoising-based upper bound (eq. 16) or a cheap stochastic estimator of it (eq. 18). These evidently similar approaches are equivalent if and only if the model is consistent across noise levels (i.e., satisfies the diffusion equation). In practice, the choice of estimation method can lead to variations of up to 0.5 bits/dimension (Kong et al., 2023), see in particular Karczewski et al. (2025) for a careful study). These approaches also apply to our model, in addition to the direct evaluation of $U_{\theta}(x,t=0)$. While this latter estimate does not correspond to a generative model (like the ODE) or a variational bound (like the SDE), its main advantage is its computational efficiency, as it does not require integrating over noise levels and gives deterministic values without averaging over noise realizations (for a divergence term or mean squared error). We note the related work of Yadin et al. (2024) which also trains an energy model with score matching and a second objective. The main conceptual difference with our approach is that we use a continuous (as opposed to discrete) time variable $t$, which offerws robustness to noise level scheduling issues and to relies on the inductive biases of existing diffusion architectures (Section 2.3).
@@ A.3 On the theoretical justification of dual score matching
A limitation of our approach is that we offer no theoretical guarantees that minimizing our objective leads to a good approximation of the energy. It would be desirable to show that our training objective (assuming infinite training data) quantitatively controls the distance between the learned energy and the data energy. This could be achieved by showing that the joint distribution $p(y,t)$ (with $\log t$ uniformly distributed in $[\log t_{\min},\log t_{\max}]$) satisfies a Poincaré inequality, i.e.,
@@
for some constant $C$ that is not too large. While we conjecture that such a result holds (or a variant, such as when considering the distribution of $(y/\sqrt{t},\log t)$, to match our noise level weighting), note that this is weaker than a control in the Kullback-Leibler divergence. Replacing $\mathrm{Var}\mathopen{}\mathclose{{}\left[U_{\theta}(y,t)-U(y,t)}\right]$ with $\operatorname{KL}\mathopen{}\mathclose{{}\left(p\,\middle\|\,p_{\theta}}\right)$ in eq. 19 would require that the model distribution $p_{\theta}(y,t)$ satisfies a log-Sobolev inequality, which could only be enforced with specific architectures. As a result, there is no control of the learned energy outside the support of the data distribution, so that out-of-distribution detection may be unreliable. It also implies our model may not be exactly normalized. Our NLL calculations are thus only approximate (although computationally cheap). These limitations are common to all current score-based and unnormalized energy-based approaches, including the related work of Yadin et al. (2024).
@@ A.4 On frequency estimation with density models
Here, we describe a counter-intuitive property of density models in high dimensions which poses a challenge for estimating frequencies. We also refer the interested reader to Theis et al. (2016) which offers related observations.
@@
A concrete example of this phenomenon can be seen in the mixture of two Gaussian distributions considered in Figure 1, where the two classes correspond to the two spheres of radius $\sqrt{d}\sigma_{i}$ for $i\in\{1,2\}$. The typical values of the energy $U_{i}=\frac{d}{2}\log\mathopen{}\mathclose{{}\left(2\pi\mathrm{e}\sigma_{i}^{2}}\right)-\log f_{i}$ are dominated by the entropy of the corresponding Gaussian. The energy is then determined by the value of the standard deviation $\sigma_{i}$ much more than the frequency $f_{i}$.
The implications of this observation for high-dimensional density models are twofold. First, estimated densities are dominated by volumes of typical sets (entropies of the mixture components), more than frequencies of different categories (one-dimensional marginals). The latter is easily learned from data, while estimating the former is a much more challenging task. Second, estimating energy up to a small relative error is not sufficient to capture observed frequencies. For our ImageNet64 model ($d=4096$), a relative error of $\sim 0.1$ dB/dimension (as estimated from the generalization experiment in Figure 2) is negligible compared to typical energy differences of $\sim 10$ dB/dimension, but the required precision to estimate frequencies with a relative accuracy of $10\%$ is $1/d=2\times{10}^{-4}$ dB/dimension.

### v3

@@ 4 Discussion
We have developed a novel framework for estimating the log probability (energy) of a distribution of images from observed samples. The estimation method is simple and robust, and converges to a stable solution with relatively small amounts of data (in our examples, $100,000$ images of size $64\times 64$ suffices). The framework leverages the tremendous power of generative diffusion models, relying on networks trained to estimate the score of the data distribution by minimizing a denoising objective. We augment this with a secondary objective that ensures consistency of the model across noise levels, and an architecture constructed from an existing denoiser so as to preserve its inductive biases. We validate our model by verifying that it achieves denoising performance equal to or surpassing the original denoiser, and NLL values comparable to the state of the art. We note that it is straightforward to extend our approach to conditional models: given conditioning information, the network needs only to compute the conditional energy function. The expression of losses $\ell_{\rm DSM}$ and $\ell_{\rm TSM}$ would be unchanged, as well as the energy architecture.
We mention two important limitations of our work. First, our energy model has a roughly quadrupled training time compared to the base score network, due to the use of double back-propagation. We believe this can be improved through architecture improvements, specialized auto-differentiation functions, or through the use of sliced score matching (Song et al., 2020). But in any case, the additional training time is offset by the dramatic efficiency gains in computing energy values. Another limitation of our approach is that we cannot provide any theoretical guarantees that minimizing our objective leads to a good approximation of the true energy. We expand on this question in Section˜A.3.
Our learned energy model provides new insights into the geometrical properties of the space of photographic images. Notably, we observe a lack of concentration of measure, with log probability values varying over a wide range and following a Gumbel distribution. As a limit distribution for the maximum, rather than the sum, of i.i.d. random variables, we expect that this might arise in a broad class of high-dimensional distributions. We also introduced a novel image- and scale-dependent measure of effective dimensionality which unifies two separate points of view developed in the literature. We demonstrate that different image neighborhoods display both high and low dimensionality over a wide range of scales. These results challenge simple interpretations of the manifold hypothesis. Moreover, the estimated dimensionalities remain too high to explain how the curse of dimensionality is avoided. This suggests that the energy landscape of natural images has additional undocumented geometrical regularity.
@@ A.1 On comparison of NLL values
Finally, a third point is that depending on the method, the NLL should be interpreted differently. Classically, $p_{\theta}$ is a normalized probability distribution, typically obtained through a change-of-variable formula (as in flow-based models), so that up to an unknown additive constant (the entropy of the data), the NLL directly evaluates the KL divergence $\operatorname{KL}\mathopen{}\mathclose{{\left(p\,\middle\|\,p_{\theta}}}\right)$. For some other models (such as VAEs), the NLL is not directly tied to a probabilistic model and is instead evaluated through a variational lower bound, leading to upper bounds on $\log p(x)$ for each $x$. The NLL however remains an upper bound on the entropy of the data. The MSE-based formulas of Kong et al. (2023) also fall in this category when using a (necessarily suboptimal) network denoiser. Lastly, approximately normalized energy-based models such as ours or CDM (Yadin et al., 2024) compute estimates of $\log p(x)$ that are neither lower nor upper bounds, so that NLL can only be interpreted as an estimate of the entropy of the data.
@@
When evaluating previously published NLL results, we found that these details were often not provided, and we thus had to make assumptions in compiling Table˜2. The “anti-aliasing” column refers to the version of ImageNet64 used, we assumed that articles that cite Van den Oord et al. (2016) for the dataset made use of the aliased version. We assume no data augmentation if none is mentioned (for TarFlow (Zhai et al., 2024), the authors mention performing center crops of the data, but an inspection of their code indicates that it has been disabled for ImageNet64). The “discreteness” column refers to the nature of the probability model and the potential conversion from continuous to discrete probabilities (“discrete” for discrete probability models, “continuous” for an additive shift of $8$ bits, and “uniform” for uniform dequantization by adding uniform noise). We assume a continuous model if no dequantization is mentioned (FM (Lipman et al., 2023) and NFDM (Bartosh et al., 2024) use a slightly different notion of uniform dequantization which improves NLL by 0.04 bits/dimension). The “type” column refers to the nature of the LLM estimate (“normalized” for an exactly normalized probability model coming from a flow-based model, “upper bound” for variational lower bounds on log probability, and “estimate” for other approaches). Finally, we indicate those methods that use a single neural function evaluation (NFE) to compute log probability in the “single NFE” column.
@@ A.2 Background on density computation in diffusion models
$\frac{\mathrm{d}x_{t}}{\mathrm{d}t}=-\frac{1}{2}\nabla_{y}U_{\theta}(x_{t},t),$ (13)
@@
$\log p_{\rm ODE}(x)=\frac{{\mathopen{}\mathclose{{\left\lVert x_{t_{\max}}}}\right\rVert}^{2}}{2t_{\max}}+\frac{d}{2}\log\mathopen{}\mathclose{{\left(2\pi t_{\max}}}\right)-\frac{1}{2}\int_{0}^{t_{\max}}\Delta_{y}U_{\theta}(x_{t},t)\mathrm{d}t.$ (14)
Note that eq.˜14 is also valid for arbitrary test points $x$, in which case the ODE (13) needs to be solved forward in time from $x_{0}=x$ at $t=0$ to $t=t_{\max}$. Equation˜14 requires estimating the divergence of the score (the Laplacian of the energy), which is typically approximated with the Hutchinson trace estimator (Hutchinson, 1989).
@@
$\log p(x)=\mathbb{E}_{y}\mathopen{}\mathclose{{\left[U(y,t_{\max})\,\middle|\,x}}\right]-\int_{0}^{t_{\max}}\mathopen{}\mathclose{{\left(td-\mathbb{E}_{y}\mathopen{}\mathclose{{\left[{\mathopen{}\mathclose{{\left\lVert x-\mathbb{E}_{x}\mathopen{}\mathclose{{\left[x\,\middle|\,y}}\right]}}\right\rVert}^{2}\,\middle|\,x}}\right]}}\right)\frac{\mathrm{d}t}{2t^{2}}.$ (15)
The first term is the effective energy at $t=t_{\max}$, which is equivalent to $\frac{d}{2}\log\mathopen{}\mathclose{{\left(2\pi\mathrm{e}t_{\max}}}\right)$ as $t_{\max}\to\infty$. The second term features the optimal mean squared error over noise levels $t\in[0,t_{\max}]$. Note that the integrand can be rewritten as $(d-d_{\rm eff}(x,t))\mathrm{d}\mathopen{}\mathclose{{\left(\frac{1}{2}\log t}}\right)$, and eq.˜15 can thus be derived by integrating the two equivalent definitions of effective dimensionality (11) and (12) (see Section˜B.3). Equation˜15 naturally leads to an upper-bound on the negative log probability of the data when replacing the optimal denoiser $\mathbb{E}_{x}\mathopen{}\mathclose{{\left[x\,\middle|\,y,t}}\right]$ with the denoiser derived from the Miyasawa-Tweedie expression, $y-t\nabla_{y}U_{\theta}(y,t)$:
$\log p_{\rm MSE}(x)=\frac{d}{2}\log\mathopen{}\mathclose{{\left(2\pi\mathrm{e}t_{\max}}}\right)-\int_{0}^{t_{\max}}\mathopen{}\mathclose{{\left(td-\mathbb{E}_{y}\mathopen{}\mathclose{{\left[{\mathopen{}\mathclose{{\left\lVert x-y+t\nabla_{y}U_{\theta}(y,t)}}\right\rVert}^{2}\,\middle|\,x}}\right]}}\right)\frac{\mathrm{d}t}{2t^{2}}.$ (16)
@@
$\log p(x)=U(x_{t_{\max}},t_{\max})-\int_{0}^{t_{\max}}\mathopen{}\mathclose{{\left(\mathopen{}\mathclose{{\left\langle\nabla_{y}U(x_{t},t),\mathrm{d}x_{t}}}\right\rangle-\frac{1}{2}{\mathopen{}\mathclose{{\left\lVert\nabla_{y}U(x_{t},t)}}\right\rVert}^{2}\mathrm{d}t}}\right).$ (17)
@@
$\log p_{\rm SDE}(x)=\frac{{\mathopen{}\mathclose{{\left\lVert x_{t_{\max}}}}\right\rVert}^{2}}{2t_{\max}}+\frac{d}{2}\log\mathopen{}\mathclose{{\left(2\pi t_{\max}}}\right)-\int_{0}^{t_{\max}}\mathopen{}\mathclose{{\left(\mathopen{}\mathclose{{\left\langle\nabla_{y}U_{\theta}(x_{t},t),\mathrm{d}x_{t}}}\right\rangle-\frac{1}{2}{\mathopen{}\mathclose{{\left\lVert\nabla_{y}U_{\theta}(x_{t},t)}}\right\rVert}^{2}\mathrm{d}t}}\right).$ (18)
Karczewski et al. (2025) show that averaging over SDE trajectories started at the same $x_{0}=x$ leads to an upper bound on the NLL, in fact recovering the one in eq.˜16.
In summary, the ODE gives deterministic probabilities exactly corresponding to the corresponding generative model (eq.˜14), while for the SDE one can choose between a denoising-based upper bound (eq.˜16) or a cheap stochastic estimator of it (eq.˜18). These evidently similar approaches are equivalent if and only if the model is consistent across noise levels (i.e., satisfies the diffusion equation). In practice, the choice of estimation method can lead to variations of up to 0.5 bits/dimension (Kong et al., 2023), see in particular Karczewski et al. (2025) for a careful study). These approaches also apply to our model, in addition to the direct evaluation of $U_{\theta}(x,t=0)$. While this latter estimate does not correspond to a generative model (like the ODE) or a variational bound (like the SDE), its main advantage is its computational efficiency, as it does not require integrating over noise levels and gives deterministic values without averaging over noise realizations (for a divergence term or mean squared error).
@@ A.3 On the theoretical justification of dual score matching
A limitation of our approach is that we offer no theoretical guarantees that minimizing our objective leads to a good approximation of the energy. It would be desirable to show that our training objective (assuming infinite training data) quantitatively controls the distance between the learned energy and the data energy. This could be achieved by showing that the joint distribution $p(y,t)$ (with $\log t$ uniformly distributed in $\log t_{\min},\log t_{\max}]$) satisfies a Poincaré inequality, i.e.,
@@
for some constant $C$ that is not too large. While we conjecture that such a result holds (or a variant, such as when considering the distribution of $(y/\sqrt{t},\log t)$, to match our noise level weighting), note that this is weaker than a control in the Kullback-Leibler divergence. Replacing $\mathrm{Var}\mathopen{}\mathclose{{\left[U_{\theta}(y,t)-U(y,t)}}\right]$ with $\operatorname{KL}\mathopen{}\mathclose{{\left(p\,\middle\|\,p_{\theta}}}\right)$ in eq.˜19 would require that the model distribution $p_{\theta}(y,t)$ satisfies a log-Sobolev inequality, which could only be enforced with specific architectures. As a result, there is no control of the learned energy outside the support of the data distribution, so that out-of-distribution detection may be unreliable. It also implies our model may not be exactly normalized. Our NLL calculations are thus only approximate (although computationally cheap). These limitations are common to all current score-based and unnormalized energy-based approaches, including the related work of Yadin et al. (2024).
@@ A.4 On frequency estimation with density models
Here, we describe a counterintuitive property of density models in high dimensions which poses a challenge for estimating frequencies. We also refer the interested reader to Theis et al. (2016), which offers related observations.
@@
A concrete example of this phenomenon can be seen in the mixture of two Gaussian distributions considered in Figure˜1, where the two classes correspond to the two spheres of radius $\sqrt{d}\sigma_{i}$ for $i\in\{1,2\}$. The typical values of the energy $U_{i}=\frac{d}{2}\log\mathopen{}\mathclose{{\left(2\pi\mathrm{e}\sigma_{i}^{2}}}\right)-\log f_{i}$ are dominated by the entropy of the corresponding Gaussian.
The implications of this observation for high-dimensional density models are twofold. First, estimated densities are dominated by volumes of typical sets (entropies of the mixture components), more than frequencies of different categories (one-dimensional marginals). (Note that the latter is easily learned from data, while estimating the former is a much more challenging task.) This observation explains empirical reports that energy models may assign higher probability to out-of-distribution samples (Nalisnick et al., 2019; Karczewski et al., 2025). Second, estimating energy up to a small relative error is not sufficient to capture observed frequencies. For our ImageNet64 model ($d=4096$), a relative error of $\sim 0.1$ dB/dimension (as estimated from the generalization experiment in Figure˜2) is negligible compared to typical energy differences of $\sim 10$ dB/dimension, but the required precision to estimate frequencies with a relative accuracy of $10\%$ is $1/d=2\times{10}^{-4}$ dB/dimension.